# Autonomy-Aware Clustering: When Local Decisions Supersede Global Prescriptions

### v1

@@ Appendix D Change in $Y$ versus $\beta$ and Critical Temperatures
Here we illustrate how $Y$ changes drastically near critical $\beta_{\text{cr}}$, and remains largely unchanged between two consecutive $\beta_{\text{cr}}$. Figure 3(a) illustrates the dataset that we consider for this illustration. It contains $3200$ data points, and we divide it into $K=16$ clusters, i.e., $16$ cluster representatives $\{y_{\ell}\}$. Each data point has local autonomy, such that it honors the prescribed cluster $15$ out $16$ times, and remaining times it overrides the prescription and uniformly associates with the remaining $15$ clusters. Figure 3(b) plots $\|\Delta Y(\beta)\|$ versus $\beta$. Note that change in $Y$ remains largely zero except at $3$ instances at which critical $\beta_{\text{cr}}$ was attained. Initially all the representatives $\{y_{\ell}\}$ are coincident, i.e., all have the same feature vector values. At first $\beta_{\text{cr}}$, $4$ distinct representative feature vectors value are formed, where each unique representative feature vector value is shared by $4$ representatives. At the second $\beta_{\text{cr}}$ each of the previous $4$ unique representative vector values give rise to $2$ unique representative vector value — making a total of $8$ unique representative feature values at this point. Here, each unique representative feature vector value is shared by $2$ representatives in $\{y_{l}\}$. At the third $\beta_{\text{cr}}$ each of the previous $8$ unique representative vectors give rise to $2$ unique representative vector values — making a total of unique $16$ representative vectors. See the .mp4 file ”SubmissionWithPT.mp4” (where plot title in every frame shows $\frac{1}{\beta}$) submitted as supplementary material for a clearer understanding.
@@ Appendix E Reinforcement-based learning for Case (C1)
Case (C1) - When $\mathcal{X}$ contains a tractable number of data points, $d(x_{i},y_{k})$ is available in closed form, and $p(l|j,i)$ is independent of $y_{\ell}$}.
@@ Appendix H Simulations with Local Autonomy
We evaluate our method on the dataset shown in Fig. 1 under a controlled form of local autonomy. Specifically, each entity $i$ accepts its prescribed cluster $j$ with probability $1-\kappa$; with probability $\kappa$ it selects an alternative cluster $k\neq j$ according to
@@
Here the cost $c_{k}(j,i)=\zeta\,d(y_{j},y_{k})+\gamma\,d(x_{i},y_{k})$. We vary the parameters $\{\kappa,\gamma,\zeta,T\}$ to study their influence on clustering outcomes. The results are shown in Fig. 5. In the visualizations, the color of each entity reflects its mixture of representative assignments, computed as a linear combination of representative colors weighted by $\pi_{Y}(j\mid i)$.
When $\kappa$ is small (Fig. 5(a1)-(a2)), cluster representatives remain close to their assigned entities. As $\kappa$ increases, deviations grow (Fig. 5(a3)-(d4)): the temperature $T$ controls the randomness of alternative selections—higher $T$ pulls all representatives toward a common location (Fig. 5(c1)-(c2)), whereas lower $T$ draws each representative toward the nearest entities (Fig. 5(a1) and 5(b1)). In the high-$\kappa$ regime (Fig. 5(d1)-(d4)), small $T$ produces pronounced shifts in which representatives migrate to the closest clusters, yielding nontrivial configurations (Fig. 5(d1), 5(d3), and 5(d4)). Note how the representative on top of a cluster is not of the same colour as the data points in that cluster — highlighting the non-trivial configurations under high $\kappa=0.9$. The parameters $\gamma$ and $\zeta$ further modulate assignment patterns, altering how entities distribute across representatives.

### v3

@@ Appendix D Change in $Y$ versus $\beta$ and Critical Temperatures
Here we illustrate how $Y$ changes drastically near critical $\beta_{\text{cr}}$, and remains largely unchanged between two consecutive $\beta_{\text{cr}}$. Figure 3(a) illustrates the dataset that we consider for this illustration. It contains $3200$ data points, and we divide it into $K=16$ clusters, i.e., $16$ cluster representatives $\{y_{\ell}\}$. Each data point has local autonomy, such that it honors the prescribed cluster $15$ out $16$ times, and remaining times it overrides the prescription and uniformly associates with the remaining $15$ clusters. Figure 3(b) plots $\|\Delta Y(\beta)\|$ versus $\beta$. Note that change in $Y$ remains largely zero except at $3$ instances at which critical $\beta_{\text{cr}}$ was attained. Initially all the representatives $\{y_{\ell}\}$ are coincident, i.e., all have the same feature vector values. At first $\beta_{\text{cr}}$, $4$ distinct representative feature vectors value are formed, where each unique representative feature vector value is shared by $4$ representatives. At the second $\beta_{\text{cr}}$ each of the previous $4$ unique representative vector values give rise to $2$ unique representative vector value — making a total of $8$ unique representative feature values at this point. Here, each unique representative feature vector value is shared by $2$ representatives in $\{y_{\ell}\}$. At the third $\beta_{\text{cr}}$ each of the previous $8$ unique representative vectors give rise to $2$ unique representative vector values — making a total of unique $16$ representative vectors. See the .mp4 file ”SubmissionWithPT.mp4” (where plot title in every frame shows $\frac{1}{\beta}$) submitted as supplementary material for a clearer understanding.
@@ Appendix E Reinforcement-based learning for Case (C1)
Case (C1) - When $\mathcal{X}$ contains a tractable number of data points, $d(x_{i},y_{k})$ is available in closed form, and $p(\ell|j,i)$ is independent of $y_{\ell}$}.
@@ Appendix H Simulations with Local Autonomy
We evaluate our method on the dataset shown in Figure 1 under a controlled form of local autonomy. Specifically, each entity $i$ accepts its prescribed cluster $j$ with probability $1-\kappa$; with probability $\kappa$ it selects an alternative cluster $k\neq j$ according to
@@
Here the cost $c_{k}(j,i)=\zeta\,d(y_{j},y_{k})+\gamma\,d(x_{i},y_{k})$. We vary the parameters $\{\kappa,\gamma,\zeta,T\}$ to study their influence on clustering outcomes. The results are shown in Figure 5. In the visualizations, the color of each entity reflects its mixture of representative assignments, computed as a linear combination of representative colors weighted by $\pi_{Y}(j\mid i)$.
When $\kappa$ is small (Figures 5(a1)-(a2)), cluster representatives remain close to their assigned entities. As $\kappa$ increases, deviations grow (Figures 5(a3)-(d4)): the temperature $T$ controls the randomness of alternative selections—higher $T$ pulls all representatives toward a common location (Figures 5(c1)-(c2)), whereas lower $T$ draws each representative toward the nearest entities (Figures 5(a1) and 5(b1)). In the high-$\kappa$ regime (Figures 5(d1)-(d4)), small $T$ produces pronounced shifts in which representatives migrate to the closest clusters, yielding nontrivial configurations (Figures 5(d1), 5(d3), and 5(d4)). Note how the representative on top of a cluster is not of the same color as the data points in that cluster — highlighting the non-trivial configurations under high $\kappa=0.9$. The parameters $\gamma$ and $\zeta$ further modulate assignment patterns, altering how entities distribute across representatives.